OpenStudy (anonymous):

Integrate ((x-1)/(x^5))^(1/2) dx show work please

OpenStudy (agreene):

\[\int\sqrt{\frac{x-1}{x^5}}dx\] is this the question? this looks like an intense problem if it is correct...

OpenStudy (anonymous):

yes

OpenStudy (agreene):

have you learned trigonometric transformation?

OpenStudy (anonymous):

No, though I don't think this problem is that complicated....I have an answer written out in the answers section of the book (Thomas Calculus) but Im still a little confused as to the mechanic of solving.

OpenStudy (agreene):

I can only think of one way of doing this, and I dont know that it would actually work.

OpenStudy (anonymous):

This problem solved using substitution.

OpenStudy (anonymous):

Hold on a second, will post a screenshot of what the book says.

OpenStudy (agreene):

yeah... i know i sub'd then trig sub'd in my head--and then it go to where im not sure if it would work.

OpenStudy (anonymous):

check that out

OpenStudy (agreene):

that solution makes no sense at all.

OpenStudy (agreene):

well... i guess it makes sense for x>0

OpenStudy (agreene):

\[\sqrt{\frac{x-1}{x^5}}=\frac1{x^2}\sqrt{\frac{x-1}{x}}\] for x>0

OpenStudy (anonymous):

What confuses me is where they would get that U from...

OpenStudy (anonymous):

would it make sense to set the entire equation inside the square root to U?

OpenStudy (agreene):

well, when you are doing those substitutions you pick your own u and find its du. They defined their u as being whats under the radical towards the end. I have a problem with the algebra, after step 2 it's pretty straightforward and correct.

OpenStudy (zarkon):

what don't you like about the algebra?

OpenStudy (agreene):

what I said above, it only holds for positive values of x.

OpenStudy (zarkon):

really?

OpenStudy (agreene):

really.

OpenStudy (zarkon):

I get then being equal for negative values

OpenStudy (zarkon):

them

OpenStudy (zarkon):

there is nothing wrong with the above solution

OpenStudy (anonymous):

Zarkon, could you write out your work please?

OpenStudy (zarkon):

you have the work

OpenStudy (anonymous):

Where does the u = 1-1/x come from?

OpenStudy (zarkon):

that is what they picked for u...since du will get rid of the \(1/x^2\)

OpenStudy (agreene):

a more basic version would be to say: \[\sqrt{x^5}=x^{5/2}\] Let's take x=-6 \[\sqrt{-6^5}=-6^{5/2}\] \[36i\sqrt{6}=-36\sqrt{6}\] one is real, the other isnt--this is a false simplification.

OpenStudy (zarkon):

\[\sqrt{x^5}=\sqrt{x^4x}=x^2\sqrt{x}\]

OpenStudy (zarkon):

again...what they did was correct

OpenStudy (zarkon):

try graphing both functions

OpenStudy (agreene):

Wolfram seems to think this is true only for x>0 as well. I think perhaps this is coming back to parenthetical insertions. That is--if you replace x with a number in parenthesis it works. http://www.wolframalpha.com/input/?i=%28%28x-1%29%2F%28x%5E5%29%29%5E%281%2F2%29+%3D+1%2Fx%5E2+%28%28x-1%29%2Fx%29%5E%281%2F2%29+

OpenStudy (jamesj):

the point is \( \sqrt{x^4} = x^2 \) for all x.

OpenStudy (agreene):

@JamesJ only with parenthetical insertions (which I was taught not to do...)

OpenStudy (zarkon):

let x=-1 \[\sqrt{\frac{x-1}{x^5}}\] \[\sqrt{\frac{-1-1}{(-1)^5}}=\sqrt{-2/(-1)}=\sqrt{2}\] \[\frac1{x^2}\sqrt{\frac{x-1}{x}}\] \[\frac1{(-1)^2}\sqrt{\frac{-1-1}{-1}}=1\cdot \sqrt{-2/(-1)}=\sqrt{2}\]

OpenStudy (jamesj):

I have no idea what parenthetical insertions are, but I imagine that if a bunch of math geeks got drunk, it would be very funny.

OpenStudy (zarkon):

I wish I could talk more ... but I'm going out to eat...I'll be back though ;)

OpenStudy (agreene):

it's what Zarkon just showed. for instance with (x^4)^(1/2)=x^2 If we insert our negatives in parenthesis it works: x=(-4) ((-4)^4)^(1/2) = (-4)^2 256^(1/2)=16 16=16 However, if we don't ..because you are adding a new, higher order, step to evaluate when you do: x=-4 (-4^4)^(1/2) = -4^2 (-256)^(1/2) = -16 16 i = -16

OpenStudy (jamesj):

From "(-4^4)^(1/2) = -4^2" to "(-256)^(1/2) = -16" is false. because, for instance, you're evaluating (-4)^2, which is 16, not -16; or (-4)^4 = 256, not -256

OpenStudy (agreene):

i dunno maybe after complex analysis my brain just became fried when it came to exponents and such

OpenStudy (agreene):

^^ James thats getting back to exactly what I was talking about. (-4)^2 =/= -4^2 Order of operations ;)

OpenStudy (jamesj):

It's not even a question of order of operations. In any case, it is true for all real numbers x that \[ \sqrt{x^4} = x^2 \] and the reworking of the integrand is correct.

OpenStudy (agreene):

So, are you saying that (-4)^2 = -4^2 ? Really?

OpenStudy (jamesj):

No. Consider the following expression \[ \sqrt{(-x)^4} \] We can write -x as =(-1)x. Hence \[ \sqrt{(-x)^4} = \sqrt{ (-1)^4x^4} = \sqrt{ 1.x^4} = \sqrt {x^4} \] Therefore \( \sqrt{(-x)^4} = \sqrt{x^4} \) and that expression is equal to \( x^2 \) for all real numbers \( x \).

OpenStudy (jamesj):

Anyway Frank, do you have your answer or not?

OpenStudy (zarkon):

what a crazy thread

OpenStudy (jamesj):

really.